I don't know how to figure this out. "My little brother left 2 minutes ago on his bike to ride with his friends to the cemetery. If he pedals at 5 mph to the cemetery, which is 1 mile from here, and I go the back way (2.5 miles) so he doesn't see the car, how fast do I have to drive Mom's car to get there at least a full minute ahead of him so I can hide and scare him?
I know the brother needs 12 minutes to go a mile, and he left 2 minutes ago, so he has 10 minutes to go.
I know the car driver needs to be there in 9 minutes. That makes 3.6 minutes a mile. What do I do from here?
Thank you!

Answer by stanbon(75887)

My little brother left 2 minutes ago on his bike to ride with his friends to the cemetery. If he pedals at 5 mph to the cemetery, which is 1 mile from here, and I go the back way (2.5 miles) so he doesn't see the car, how fast do I have to drive Mom's car to get there at least a full minute ahead of him so I can hide and scare him?
------------------
Brother DATA:
rate = 5 mph ; distance = 1 miles ; time = d/r = 1/5 = 12 minutes
--------------------------
You DATA:
distance = 2.5 miles ; time = 11 minutes ; rate = d/t = 2.5/11 miles/min
---
11 min = 11/60 hr = 0.1833 hr
----
2.5/11 (miles/min) = 2.5/0.1833 = 13.64 miles/hr
=================================
Cheers,
Stan H.
================